An object is thrown downward from the top of a tower. The distance in feet that it falls in t seconds is given by the following.
s(t) = 16t2 + 14t
(c) Find the average velocity from t = 3 seconds to t = 5 seconds.
3 ft/sec

Answer by nerdybill(7384)

s(t) = 16t2 + 14t
average velocity from t = 3 seconds to t = 5 seconds
.
s(3) = 16*3^3 + 14(3)
s(3) = 16*27 + 14(3)
s(3) = 474
.
s(5) = 16*5^3 + 14(5)
s(5) = 16*125 + 14(5)
s(5) = 2070
.
Avg vel = (s(5)-s(3))/(5-2)
Avg vel = (2070-474)/(3)
Avg vel = (1596)/(3)
Avg vel = 532 feet/sec